OpenStudy (anonymous):

if y varies directly with x and y = 4 when x = 8, then find y when x = 1. a. y = -3/5 b. y = 1/2 c. y = -1/2 d. y = -7/10

jhonyy9 (jhonyy9):

4---8 y---1 -------- y=4/8 =1/2

jhonyy9 (jhonyy9):

b. is right

OpenStudy (anonymous):

Yeah, I agree. Draw a graph- the slope is 2. And that's (almost) my username!